My math question is as follows: Find two consecutive odd integers that twice the larger is 15 more than 3 times the smaller.

Answer by stanbon(75887)

Find two consecutive odd integers that twice the larger is 15 more than 3 times the smaller.
-------
1st: 2x-1
2nd: 2x+1
---------------
Equation:
2(2x+1) = 3(2x-1)+15
-----
4x + 2 = 6x + 12
2x = -10
x = -5
----
1st: 2x-1 = -11
2nd: 2x+1 = -9
=====================
Cheers,
Stan H.
=======================